OpenStudy (anonymous):

|2n - 3| = 5 can u explain how to solve and help me solve? I dont really get what the lines are for either

OpenStudy (nincompoop):

it is a piecewise-defined where n ≤ 0 and n > 0

OpenStudy (anonymous):

so how do u solve

OpenStudy (nincompoop):

solve two parts 1*(2n - 3) = 5 -1*(2n - 3) = 5

OpenStudy (anonymous):

i dont get it

OpenStudy (nincompoop):

can you solve those two I've written out?

OpenStudy (anonymous):

no... if i cant solve the problem i began with than i cant solve the one u wrote... sorry math is my worst subject

OpenStudy (anonymous):

im so confused

OpenStudy (nincompoop):

you don't know distributive property?

OpenStudy (anonymous):

i do

OpenStudy (nincompoop):

THEN SOLVE those two I've given you

OpenStudy (anonymous):

its the same but negative

OpenStudy (nincompoop):

I want to see you solve it

OpenStudy (nincompoop):

when someone is being patient teaching you, pay attention and heed to request

OpenStudy (anonymous):

i dont get it!

OpenStudy (nincompoop):

then you don't know distributive property

OpenStudy (anonymous):

i do tho

OpenStudy (nincompoop):

then why can't you solve what I wrote?

OpenStudy (anonymous):

because i dont get it... how hard is it to understand what im saying

OpenStudy (nincompoop):

solve two parts 1*(2n - 3) = 5 -1*(2n - 3) = 5 if you don't solve those two, I am done with you

OpenStudy (anonymous):

math is my worst subject

OpenStudy (anonymous):

i dont get how to solve it

OpenStudy (anonymous):

2n-3=5

OpenStudy (nincompoop):

so you DO NOT KNOW DISTRIBUTIVE PROPERTY

OpenStudy (anonymous):

-2n- -3= 5

OpenStudy (anonymous):

there i did it but its wrong isnt it

OpenStudy (nincompoop):

did you solve for n? SOLVE FOR n in each case

OpenStudy (nincompoop):

yeah it is wrong because when you are multiplying two negatives, you get positive

OpenStudy (anonymous):

how is it wrong i dont have any negative numbers only the 1 that that u gave me

OpenStudy (nincompoop):

-1*(2n - 3) = 5 THERE

OpenStudy (anonymous):

-2n - -3 = 5

OpenStudy (nincompoop):

what?

OpenStudy (nincompoop):

what is -1 * -3?

OpenStudy (anonymous):

its not negative three its minus 3

OpenStudy (nincompoop):

minus 3 is THE SAME AS NEGATIVE 3

OpenStudy (anonymous):

oh okay so than -2n-3=5

OpenStudy (anonymous):

-2n=8

OpenStudy (anonymous):

n is -4

OpenStudy (nincompoop):

?

OpenStudy (anonymous):

am i rright

OpenStudy (nincompoop):

learn arithmetic

OpenStudy (anonymous):

these are the possible answers {1, 4} {-1, 4} {1, -4} {-1, 1}

OpenStudy (nincompoop):

I am done with you

OpenStudy (anonymous):

wait!

OpenStudy (anonymous):

why cant u tell me if i am right or wrong

OpenStudy (nincompoop):

I already told you that you were wrong

OpenStudy (anonymous):

well i gave u all the options for the answers... which one is it

OpenStudy (anonymous):

i thought the 3rd

OpenStudy (anonymous):

and plus i dont need arithmetic this is 4th grade ap math

OpenStudy (nincompoop):

I think you need to hone skill in arithmetic, buddy